Find a quadratic equation for the solutions of 1/3 and 5


Answer by ewatrrr(24785)

 
Hi,
Find a quadratic equation for the solutions of 1/3 and 5
(x-1/3)(x-5) = 0
F First terms
O Outside terms
I Inside terms
L Last terms
x^2 - 1/3x - 5x + 5/3 = 0
x^2 - (16/3)x + 5/3 = 0
3x^2 -16x + 5 = 0
y = 3x^2 -16x + 5